On July 1, 2017, Crane Company purchased factory equipment for $293000. Salvage value was estimated to be $8500. The equipment will be depreciated over five years using the double-declining balance method. Counting the year of acquisition as one-half year, Crane should record depreciation expense for 2018 on this equipment of

$70320.
$93760.
$91200.
$117200.
Answer #1

Double Declining Depreciation = Beginning Book Value x 2 times straight line rate
straight line rate = 1/5 x 100 = 20%

Depreciation for 2017 = $293000 x 40% x 6/12 = $58600
Depreciation for 2018 = ($293000 - 58600) x 40% = $93760

Answer is b. $93760